The ratio of chairs to tables is 7:8. If there are 14 chairs, how many tables are there?
1
Expert's answer
2019-06-16T17:09:03-0400

If the ratio is 7:8,we can rewrite it like 7x:8x (7x is the number of chairs, 8x is the number of tables), that is the ratio of chairs to tables, then we get equation 14=7x,divide both sides by 7,get x=2,then the number of tables equals 8*x=8*2=16 tables.

Check the answer.The ratio of chairs to tables is 14:16 which is the same as 7:8.